OpenStudy (anonymous):

What is the equation of a circle with center (-3, -1) that contains the point (1, 2)? A. (x + 3)2 + (y + 1)2 = 25 B. (x − 3)2 + (y − 1)2 = 36 C. (x + 3)2 + (y − 1)2 = 36 D. (x − 1)2 + (y − 2)2 = 25

OpenStudy (anonymous):

\[equation\ of \ circle \ with \ center \ (a,b) \ radius\ r \ \ and\ general \ points \ (x,y) \\ on \ circumference \ = (x-a)^2 + (y-b)^2= r^2 \\ here \ (a,b) = (-3,-1) \\ r^2= (-3-1)^2+(-1-2)^2=25\\therefore \ equation \ of \ circle \ is \ [x-(-3)]^2+[y-(-1)]^2=25 \\ [x+3]^2+[y+1]^2=25\] Hence option A is correct.

OpenStudy (anonymous):

figured that was the answer, ^-^ thanks